

********************************************************************
UARS WEEKLY STATUS REPORT - 09 Oct 1998
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM successfully gathered its normal solar, offset, and occulation
data during this past week without apparent problems.
***********
PEM
***********
PEM is now conducting tests of the OBC activation. PEM activation
occurs during daylight only, when there is adequate power generated from
the UARS solar array. Currently, the PEM magnetometer and medium energy
spectrometers (particles below 30 keV) activate correctly under OBC control.
Tests conducted on the high energy particle spectrometers failed to enable the
instrument correctly. These tests show failure to enable high voltage power,
however it is felt that the communication rate is currently too high to the HEPS
microprocessor/PEM processor bus and this effect is masking high voltage
activity. The PEM team is investigating this result. High voltage tests are
planned with out processing the majority of the microprocessor commands, which
reset detector characteristics for scientific validity. If the high voltage
tests are successful, then additional tests of the communication rate will be
needed to configure the microprocessor to the proper state.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
MLS was reactivated successfully on Oct 6 and is scheduled to operate though
Oct 10. The MLS team is operating the instrument in a mode for diagnosing the
reoccurance of slips in the antenna scan mechanism, and to attempt to find a
method of operating the instrument which minimizes the impact of these slips on
scientific data.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
We welcome Monica Todirita as a new member of the Mission
Planning Group.
***********
CDHF
***********
CDHF personnel will be closely monitoring the temperature and humidity
this upcoming weekend (10/09/98 - 10/12/98) due to the planned chilled
water and steam outage. If needed, Building 23 will have portable A/C units
available for use. Current plans do not include any interruption in
CDHF services during this time.
The CDHF has received a total of 425 CD-ROM requests as of 09-October-1998.
There were three orders received and shipped.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
78 days HAL (4.80) PROD (L1) (1873R-1638C-24F=211L)
135 days HAL (4.80) PROD (L3) (1873R-1017C=856L)
62 days SOL (4.62) PROD (L2-L3) (2349R-1748C=601L)
4 days SUS (7.00) TEST (4R-4C=0L)
CDHF pending software upgrades: IDL 5.1, OpenIngres.
***********
FOT
***********
REPORTING PERIOD
This report covers 26 September 1998 (Orbit 38474, GMT Day 269) through 02
October 1998 (Orbit 38578, GMT Day 275).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 269/2321 38489 Off
HRDI 272/1950 38532 Day Only
PEM 273/2004 38547 Testing
PEM 274/1517 38559 Testing
PEM 274/2010 38562 Testing
PEM 275/1214 38573 Testing
A summary of Instrument operations supplied by the Mission Planning Group
for the month of September 1998 is contained in Appendix F, UARS Instrument
Operations Summary.
The beta angle ranged between 17.1 degrees to 10.1 degrees for this report
period. The beta angle is now increasing toward a maximum angle of 48.0
degrees on 17 October 1998 (DOY 290).
A Reverse-to-Forward Yaw Around maneuver (number 72) was successfully
executed in Orbit 38529 beginning 272/14:24:59 and completing at
272/15:02:09 on 29 September 1998. Due to the SA position at 269 degrees,
the effective night length of the spacecraft after the maneuver was
approximately 50 minutes. Post maneuver analysis of data indicate that
during the Yaw Around, the end-of-night load bus voltage was 24.0 volts and
the average depths of discharge (DODs) on Batteries 2 and 3 were
approximately 23.0 and 23.0 percent respectively.
A manually generated SSPP stellar offset experiment requested by SOLSTICE
was performed on 29 September 1998 (GMT Day 272).
Initial operation of HRDI utilizing the Power Monitor "day only"
operational capability occurred on 29 September 1998. Day Only HRDI
operation has changed the heater control requirements for the instrument.
HRDI heater cycling is under analysis. PEM transition to day-only
operation was initiated on 30 September and continued through 02 October
1998. PEM operation is partial and investigation is underway to verify
command sequences for high voltage operation.
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 0.9 to 1.3 degrees C. Additional battery performance data
for each battery on 25 September 1998 and 02 October 1998 are provided in
Appendix D.
The clock error ranged from 0.3 to 9.9 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = 01
% OUT-OF-LIMITS = 08
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 06
% HALOE 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 04
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 98/210 - 2287.496941
Transponder B: 98/211 - 2287.499161
GROUND SYSTEM OPERATION
Three (3) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 98/274 show the observatory data loss to be
21 hours, 08 minutes, 31.876 seconds (an increase of 1 minute and 45.536
seconds since last report period). This is a 0.0343 percent data loss
which equals a 99.9657 percent data capture for the mission.
The increase of 1 minute and 45.536 seconds occurred on Day 270 and was due
to hits in the data, line errors, and insufficient overlap.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
FOT continues to improve a power profile for UARS based on the current
status of the Modular Power System.
Investigation continues into possible impacts of the Leonid meteorite
shower in November 1998.
YEAR 2000 OPERABILITY SUPPORT
UARS Test and Training Simulator (UTTS) - No change.
User Planning System (UPS) - The FOT was unsuccessful in transferring
simulated Y2K TDRS contacts to the CMS.
Multi Satellite Operations Control Center (MSOCC) - The FOT generated some
sample ATC, RTS, and ephemeris loads for Y2K POCC/AP testing.
Unfortunately, the FOT is currently unable to transfer and block these Y2K
loads using the back-up DFLX system (DFLX2). This is due to the inability
in getting the loads to DFLX2. The original testing plan restricts the FOT
from using the operational DFLX for Y2K testing, so until this issue is
resolved, the Y2K CMS loads cannot be processed. The Y2K CMS software
developer is currently working on resolving the problems between UCMSB and
DFLX2. In the meantime, permission to use the prime CMS and DFLX to
transfer Y2K loads is being investigated
At the POCC/AP Year 2K test team weekly meeting, the FOT provided Mr. Brown
with a draft memo which answered the majority of the action items from last
weekUs Y2K meeting. Also, at this meeting, Mr. Brown showed the FOT
several printouts indicating that a previously identified Y2K problem
(regarding invalid year depiction on the Event and other reports) had been
resolved.
Command Management System (CMS) - The FOT assisted in CMS Y2K test team in
further testing on the back-up CMS system this week. The two groups
attempted to finalize as much testing as possible so the appropriate test
reports could be written to meet the 30 September test deadline. The FOT
still has yet to receive a valid S/C Orbital Events file from FDF, so
approximately 50 percent of the CMS functions could not be tested.
Unfortunately, the majority of the CMS processing accesses the CMS
Scheduling Events Pool file, which must contain UARS orbit numbers (which
are extracted from the missing S/C Orbital Events file). Although the 30
September deadline has arrived, further testing will continue as soon as a
valid Orbital Events file is provided by FDF.
The UPS has a "test" mode in which the user can transmit "non-confirmed"
TDRS events to the CMS, but CMS does not appear to be accepting the files
from the UPS and processing them correctly. This process, as well as other
alternatives to getting Y2K TDRS data on the CMS is still being
investigated.
Flight Dynamics Facility (FDF) - See Command Management System. FOT has
yet to receive a valid S/C Orbital Events file from FDF.
IP NETWORK TESTING
Space Network (SN) - Declared successful at the IP Transition ORR.
Testing with the NCC 98 system to support routine flight operations is
being planned.
Deep Space Network (DSN) and Ground Spaceflight Tracking and Data Network
(GSTDN) - No change and incomplete.
UARS PROPULSION MODULE B/D THRUSTER LINE FILL
The B/D thruster refill effort continues. The Eaton Latch Valve test
report is still being reviewed by the UARS FOT. Jackson and Tull reported
that the latch valve drivers and cables for the propulsion line refill test
should be available next week.
UARS RE-ENGINEERING - TELEMETRY PROCESSING
Flight operations is providing support to the re-engineering effort to
process the UARS telemetry on the TOMS/EP control center computer system.
The software performance continues to be evaluated.
MEETINGS
The FOT met with Eddie Brown and Rick Sanders of the POCC/AP Year 2K test
team on 30 September to discuss possible Year 2K impacts on the UARS
POCC/AP software.
UARS supported the IP Transition Operational Readiness Review (ORR)
conducted on 01 October 1998. One chart was prepared stating that the
TDRSS testing was successful but the Deep Space Network (DSN) testing is
incomplete.
The next battery telecon is scheduled for 20 October 1998 at 1:00 pm.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is scheduled for scheduled for 02
November 1998.
UARS will pass through the Leonid meteor shower on 17 November 1998 at
19:43 UTC (+/- 3 hours).
The United States Naval Observatory (USNO) issued a Time Service
announcement advising that the International Earth Rotation Service (IERS)
will retard UTC time by 1.0 second at the end of December 1998. This means
the addition of one leap second as the last second of 1998.
The next annular solar eclipse will occur on 16 February 1999 over the
Indian Ocean, Australia and the South Pacific Ocean.
PERSONNEL CHANGES
None
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
Appendix F, UARS Instrument Operations Summary
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 02 October 1998
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
38489 269/2321 98-365 HALOE Power OFF
38532 272/1950 98-366 HRDI Power On thru PMON
38547 273/2004 98-367 PEM PEM/PMON control
38559 274/1517 98-368 PEM/PMON Testing
38562 274/2010 98-369 PEM/PMON Testing
38573 275/1214 98-370 PEM/PMON PEM Test
38575 275/1643 98-371 OBC/PMON HRDI Heater
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 02 October 1998
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
WIFWTEMP Y-LO WINDII Filter Wheel Temp 38474
WIISUTEMP R-LO WINDII ISU Temperature 38474-496,38530,31
HRPSEQBIMON Y-HI PSE Quiet Bus Current Mon 38479
NBRECBI Y-HI NBTR B Total Current 38479,542
ACTEMPANA1 Y-LO Temperature AN1 38545
HRPTSPBIMON R-LO PTS Pulse Bus Current Mon 38553
PEPCUZEPSIE R-LO PCU/ZEPS Current Monitor 38553,554
PEPCUPRIIE R-LO PCU Primary Current Monitor 38575
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
WIISUTEMPRL WINDII ISU Temperature 38474-496,38530
_______________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 02 October 1998
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
None this report period
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 02 October 1998
25 Sept 1998 - GMT Day 268 Beta = 17.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-4.31 +0.0/ +0.0 18.1 0.000 0.0
2 +50.4/0.0 +6.76/+5.15 +31.6/-16.0 24.2 1.023 22.8
3 +33.6/-44.8 +5.47/+3.87 +30.4/-16.4 24.2 1.021 22.2
02 Oct 1998 - GMT Day 275 Beta = 10.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.68/-4.94 +0.0/ +0.0 18.1 0.000 0.0
2 +44.8/-5.6 +4.51/+3.55 +30.4/-13.2 26.4 1.022 20.4
3 +33.6/-22.4 +3.23/+2.60 +28.8/-12.8 26.4 1.022 19.7
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes (V/T 5)
APPENDIX E
UARS ANOMALY REPORTS
Week ending 02 October 1998
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-064 FOT 38482 269/1154 N/A Unable to Command/No Loss
98-065 FOT 38496 270/1016 20220 Data Drop-out/Yes
98-057 FOT 38521 272/0240 14170 Generic Late Acquisition
98-066 FOT 38572 275/1214 14170 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-057 FOT Sept 98 See List 14170 GENERIC LATE ACQ
98-061 FOT 38385 263/00:30 20202 AP CRASH
98-062 FOT 38447 266/04:01 TMON 1 & 2 TRIP
98-063 FOT 38456 267/18:43 ATC LOAD ERRORS
98-064 FOT 38482 269/11:54 UNABLE TO COMMAND
98-065 FOT 38496 270/10:16 20220 DATA DROP OUTS
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-066 FOT Oct 98 See List 14170 GENERIC LATE ACQ
_______________________________________________________________
APPENDIX F
UARS INSTRUMENT OPERATIONS SUMMARY
Week ending 02 October 1998
INSTRUMENT SUMMARY FOR MONTH OF September 1998 (hours)
UNPLANNED UNSCHEDULED
ON Day- OFF OFF OFF
Only (Insufficient (Other
Instrument (Actual) Ops (Scheduled) Power) Problem) COMMENTS
---------- ------ --- ----------- ----------- ----------- --------------
ACRIM 544 169 7 0
HALOE 615 99 6 0
HRDI 166 4 550 0 0 See Note 1
MLS 164 556 0 0
PEM (axis) 41 439 0 0
PEM (zeps) 97 623 0 0
PEM (neps) 0 720 0 0
SOLSTICE 713 0 7 0
SUSIM 540 170 10 0
WINDII 196 524 0 0
Note 1. Day-Only operation started on 30 September 1998
Cold side of spacecraft was facing northward: 28.5 days
Cold side of spacecraft was facing southward: 1.5 days